I have no idea how to work this problem, I keep getting the wrong answer
please help
Solve:
(y^2-5y)^2-8(y^2-5y)-84=0

Answer by scott8148(6628)

this is a quadratic in a quadratic

factoring ___ [(y^2 - 5y) - 14][(y^2 - 5y) + 6] = 0

y^2 - 5y - 14 = 0 ___ factoring ___ (y - 7)(y + 2) = 0 ___ y = 7 , y = -2

y^2 - 5y + 6 = 0 ___ factoring ___ (y - 3)(y - 2) = 0 ___ y = 3 , y = 2